OpenStudy (anonymous):

PLEASE HELP! Solution of 2 [7x-4] +3=51?

OpenStudy (karatechopper):

When trying to isolate x, what is the first step possible?

OpenStudy (anonymous):

idk im terrible at this stuff :/

OpenStudy (anonymous):

2(7x-4)+3 = 51 14x - 8 + 3 = 51 14x -5 = 51 14x = 56 x = 4

OpenStudy (tkhunny):

Why did you use [ ] instead of ( )? Were you trying to indicate Absolute Values?

OpenStudy (anonymous):

yes!!!

OpenStudy (anonymous):

help please?

OpenStudy (anonymous):

2[7x-4] = 48 [7x-4] = 24 7x-4 = -24 OR 7x-4 = 24 7x = -20 OR 7x = 20 x = -20/7 OR 7x = 20/7

OpenStudy (anonymous):

sorry, the second part should be OR 7x = 28 -> x=4